OpenStudy (anonymous):

How do you determine the direction for the normal: (3,2,4)

OpenStudy (anonymous):

Direction vector*

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Sorry. I didn't phrase my question right. I know how to find the direction vector of a line from the normal vector of the line when in 2 space since they are perpendicular to each other. For ex. The normal may be (2,1) so I could assume the director vector if line would be (-1,2) or (1,-2) but would I do so, if it's in 3 space

OpenStudy (anonymous):

Of the line*

OpenStudy (anonymous):

In 3D, there are infinity many vectors that are perpendicular to a given line. Of you can just imagine the vector rotating around the line. There is a test whether two vectors in 3D are perpendicular, and that is called the dot product. two vectors are perpendicular if the dot product between them equals 0

OpenStudy (anonymous):

suppose vector (a,b,c). (a,b,c) dot (3,2,4) = 0 2a + 3b + 4c = 0 so all vectors (a,b,c) that satisfy the equation above are perpendicular to (4,2,4)

OpenStudy (anonymous):

*(3,2,4)*

OpenStudy (anonymous):

So , -2, -3, and -4 could possibly be the direction vector of the line?

OpenStudy (anonymous):

the equation was supposed to be 3a + 2b + 4c = 0 so if (a,b,c) = (-2,-3,-4), then 3(-2) + 2(-3) + 4(-4) = -28, which is not 0. Thus (-2,-3,-4) is not perpendicular to (3,2,4)

OpenStudy (anonymous):

Thanks!